# BiSSL: Bilevel Optimization for Self-Supervised Pre-Training and Fine-Tuning

### v1

@@ 1 Introduction
In the absence of sufficient labeled data, self-supervised learning (SSL) has emerged as a promising approach for training deep learning models. Rather than relying solely on labeled data, the SSL framework aims to learn representations from unlabeled data which proves beneficial for subsequent use on various downstream tasks. These representations are learned by solving a pretext task, which utilizes supervisory signals extracted from the unlabeled data itself. Extensive efforts have been made into designing effective pretext tasks which in recent years have led to state-of-the-art or competitive performance in various fields such as computer vision (Chen et al., 2020b; Bardes et al., 2022; He et al., 2020; Grill et al., 2020; Caron et al., 2021; He et al., 2022), audio signal processing (Schneider et al., 2019; Baevski et al., 2020; Hsu et al., 2021; Niizumi et al., 2021; Chung & Glass, 2018; Chung et al., 2019; Yadav et al., 2024) and natural language processing (Devlin et al., 2019; Lewis et al., 2019; Brown et al., 2020; He et al., 2021).
@@
In this study, we propose BiSSL, a novel training framework that leverages BLO to enhance the alignment between the pretext pre-training and downstream fine-tuning stages in SSL. Acting as an intermediate training stage within the SSL pipeline, BiSSL frames the pretext and downstream task objectives as the lower- and upper-level objectives in a BLO problem - a challenging approach that has not been explored until now. The objectives in BiSSL are connected by substituting the lower-level backbone solution for the upper-level backbone parameters, while simultaneously enforcing the lower-level backbone solution to resemble the upper-level backbone parameters. This approach more explicitly captures the interdependence between pretext pre-training and downstream fine-tuning, potentially leading to a lower-level backbone better aligned with the downstream task. Figure 1 compares the conventional SSL pipeline with our suggested pipeline involving BiSSL. Additionally, we propose a training algorithm for BiSSL and demonstrate that it consistently improves or maintains comparable downstream performance across a range of image classification datasets. For our experiments, we use SimCLR (Chen et al., 2020b) to pre-train a ResNet-18 backbone (He et al., 2016) on the unlabeled partition of the STL10 dataset (Coates et al., 2011), a setup offering suitable model capacity and dataset complexity while being less resource-intensive than larger-scale alternatives. The implementation and pre-trained model weights will be made publicly available. [^1] https://github.com/ICLR25-10484/ICLR25_10484_BiSSL
@@ 3.3 Upper-level Derivative
Due to the dependence of the lower-level solution on the upper-level parameters, the first term of equation 4 includes the implicit gradient (IG) of the implicit function ${\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D})$. To simplify notation, we let $\nabla_{{\bf\it\xi}}h({\bf\it\xi})|_{{\bf\it\xi}={\bf\it\psi}}:=\nabla_{{\bf\it\xi}}h({\bf\it\psi})$ when it is clear from context which variables are differentiated with respect to. Following an approach similar to Rajeswaran et al. (2019), with details on the derivations and underlying assumptions outlined in Appendix A, the IG in equation 4 can be explicitly expressed as
@@ 3.4 Training Algorithm and Pipeline
From Appendix A, we get that ${\bf\it\theta}^{\ast}_{P}({\bf\it\theta}_{D})$ must fulfill the stationary condition $\nabla_{{\bf\it\theta}}\quantity(\mathcal{L}^{P}({\bf\it\theta},{\bf\it\phi}_{P})+\lambda r({\bf\it\theta}_{D},{\bf\it\theta}))|_{{\bf\it\theta}={\bf\it\theta}^{\ast}_{P}({\bf\it\theta}_{D})}=\mathbf{0}$ to justify the explicit expression of the implicit gradient in equation 6. This means that executing Algorithm 1 using random initializations of ${\bf\it\theta}$ and ${\bf\it\phi}_{P}$ will likely not suffice. The same applies to ${\bf\it\phi}_{D}$, as a random initialization of ${\bf\it\phi}_{D}$ typically leads to rapid initial changes of the backbone parameters ${\bf\it\theta}_{D}$ during fine-tuning. This would then likely violate the assumed stationary condition due to the dependence between ${\bf\it\theta}_{D}$ and ${\bf\it\theta}_{P}$ through the regularization objective $r$. Figure 1 illustrates the suggested pipeline alongside the conventional SSL pipeline. First conventional pretext pre-training is performed on the unlabeled dataset $\mathcal{D}^{P}$ to obtain initializations of ${\bf\it\theta}$ and ${\bf\it\phi}_{P}$. Next, the downstream head is fitted on top of the frozen backbone ${\bf\it\theta}$ using the downstream dataset $\mathcal{D}^{D}$, which provides an initialization of the downstream head parameters ${\bf\it\phi}_{D}$. Then, BiSSL training is conducted as outlined in Algorithm 1, yielding an updated configuration of backbone parameters ${\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D})$. These updated backbone parameters are subsequently used as an initialization for the final supervised fine-tuning on the downstream task.
@@ 4.2.2 BiSSL Setup
The backbone ${\bf\it\theta}$ and projection head ${\bf\it\phi}_{P}$ are initialized by self-supervised pre-training using a setup almost identical to the baseline pretext task training setup in Section 4.2.1. The only difference is that this training stage is conducted for $500$ epochs instead of $600$ epochs. This adjustment is made because the BiSSL training stage will conduct what is roughly equivalent to $100$ pretext epochs, as detailed more specifically in the composite configuration paragraph below. This ensures that the total number of pretext pre-training steps is comparable to those conducted in the baseline setup.
@@
STL10L $\mathbf{90.6\pm 0.1}$ $90.3\pm 0.1$ $\mathbf{+0.3}$ $\mathbf{99.7\pm 0.0}$ $99.6\pm 0.0$ $\mathbf{+0.1}$
Flowers $\mathbf{74.8\pm 0.4}$ $73.4\pm 0.4$ $\mathbf{+1.4}$ $90.0\pm 0.3$ $90.0\pm 0.4$ $0.0$
Cars $72.4\pm 0.3$ $72.7\pm 0.5$ $-0.3$ $91.2\pm 0.2$ $91.4\pm 0.4$ $-0.2$
Aircrafts $46.8\pm 0.8$ $46.1\pm 0.9$ $+0.7$ $78.5\pm 0.6$ $79.3\pm 0.6$ $-0.8$
DTD $\mathbf{52.1\pm 0.5}$ $49.3\pm 0.5$ $\mathbf{+2.8}$ $79.5\pm 0.3$ $79.1\pm 0.4$ $+0.4$
Pets $\mathbf{67.6\pm 0.4}$ $65.0\pm 0.5$ $\mathbf{+2.6}$ $\mathbf{91.9\pm 0.4}$ $90.7\pm 0.3$ $\mathbf{+1.2}$
FMNIST $94.3\pm 0.1$ $94.1\pm 0.1$ $+0.2$ $100.0\pm 0.0$ $100.0\pm 0.0$ $0.0$
CIFAR10 $93.8\pm 0.2$ $93.8\pm 0.1$ $0.0$ $\mathbf{99.9\pm 0.0}$ $99.8\pm 0.0$ $\mathbf{+0.1}$
CIFAR100 $73.0\pm 0.2$ $73.2\pm 0.2$ $-0.2$ $\mathbf{93.7\pm 0.1}$ $92.8\pm 0.1$ $\mathbf{+0.9}$
Caltech-101 $\mathbf{80.7\pm 0.5}$ $78.1\pm 0.5$ $\mathbf{+2.6}$ $\mathbf{95.3\pm 0.1}$ $94.7\pm 0.2$ $\mathbf{+0.6}$
Food $71.8\pm 0.2$ $71.7\pm 0.2$ $+0.1$ $90.5\pm 0.1$ $90.4\pm 0.1$ $+0.1$
SUN397 $\mathbf{41.0\pm 0.3}$ $40.0\pm 0.3$ $\mathbf{+1.0}$ $\mathbf{71.0\pm 0.4}$ $69.9\pm 0.4$ $\mathbf{+1.1}$
CUB200 $45.6\pm 0.3$ $45.7\pm 0.4$ $-0.1$ $\mathbf{75.2\pm 0.2}$ $70.7\pm 0.6$ $\mathbf{+4.5}$
VOC07 $\mathbf{*60.4\pm 0.2}$ $*58.6\pm 0.3$ $\mathbf{+1.8}$- -- --
@@ 4.3.1 Performance over varying pre-training epochs
Table 1 demonstrates that BiSSL can significantly enhance classification accuracies across various datasets. To determine whether this improvement persists across different levels of pretext pre-training, additional experiments are conducted using varying numbers of total pre-training epochs. The accuracies are compared with the total number of pretext pre-training epochs kept roughly the same, so the pretext warm-up in the BiSSL pipeline is conducted for $100$ epochs less than the baseline pretext pre-training, as justified in the ‘Composite Configuration Details of BiSSL” paragraph of Section 4.2.2. The flowers dataset, which showed substantial benefits from BiSSL on the top-1 classification accuracies, is selected for these experiments. The results presented in Figure 2 reveal that BiSSL generally sustains the relative performance gap over the baseline, regardless of the pre-training duration. This suggests that the advantages conferred by BiSSL are not contingent on the amount of pre-training. Rather, this indicates that BiSSL may provide a more efficient learning trajectory, stemming from the enhanced information sharing it facilitates between the pretext and downstream tasks.
@@ 4.4 Visual Inspection of Latent Features
Figure 3: Visualization of features from backbones trained using pretext pre-training exclusively and backbones derived from lower-level backbones obtained after applying BiSSL, respectively. Features are extracted from the test partition of the flowers dataset. Each color represents a different class. Details are outlined in Section C.1 of Appendix C.
@@ 5.1 Future Work
Formulating the self-supervised pipeline as a bilevel optimization problem can be approached through various strategies, each involving trade-offs in computational complexity and theoretical justification. While this study demonstrates a promising method for enhancing downstream performance, further exploration of alternative formulations is necessary to identify potentially more optimal setups. Moreover, although BiSSL generally applies to any downstream task and model size in theory, our experiments were confined to small-scale image classification due to resource constraints. Consequently, it remains an open question whether BiSSL can scale effectively to larger setups and other downstream tasks.
@@ Acknowledgments
The authors would like to thank Sijia Liu and Yihua Zhang (Michigan State University) for providing valuable feedback in a discussion, which helped to refine and solidify our perspective on the topic of integrating bilevel optimization in deep learning.
Appendix A Derivation of the Implicit Gradient
@@ B.5 Downstream Fine-Tuning after BiSSL
STL10L $0.0008$ $0.00641$
Flowers $0.0007$ $0.00078$
Cars $0.028$ $0.00453$
Aircrafts $0.0126$ $0.00459$
DTD $0.005$ $0.00003$
Pets $0.0011$ $0.00852$
FashionMNIST $0.0031$ $0.00983$
CIFAR10 $0.0007$ $0.00611$
CIFAR100 $0.0013$ $0.00322$
Caltech-101 $0.0004$ $0.00013$
Food $0.0202$ $0.00697$
SUN397 $0.0015$ $0.00298$
CUB200 $0.0003$ $0.00495$
VOC2007 $0.0004$ $0.00003$

### v2

@@ 1 Introduction
In the absence of sufficient labeled data, self-supervised learning (SSL) has emerged as a promising approach for training deep learning models. Rather than relying solely on labeled data, the SSL framework aims to learn representations from unlabeled data which proves beneficial for subsequent use on various downstream tasks. These representations are learned by solving a pretext task, which utilizes supervisory signals extracted from the unlabeled data itself. Extensive efforts has gone into designing effective pretext tasks, achieving state-of-the-art or competitive performance in various fields such as computer vision (Chen et al., 2020b; Bardes et al., 2022; He et al., 2020; Grill et al., 2020; Caron et al., 2020; 2021; He et al., 2022; Oquab et al., 2024), audio signal processing (Schneider et al., 2019; Baevski et al., 2020; Hsu et al., 2021; Niizumi et al., 2021; Chung & Glass, 2018; Chung et al., 2019; Yadav et al., 2024) and natural language processing (Devlin et al., 2019; Lewis et al., 2019; Brown et al., 2020; He et al., 2021; Touvron et al., 2023).
@@
In this study, we propose BiSSL, a novel training framework that leverages BLO to enhance the alignment between the pretext pre-training and downstream fine-tuning stages in SSL. Acting as an intermediate training stage within the SSL pipeline, BiSSL frames the pretext and downstream task objectives as the lower- and upper-level objectives in a BLO problem - a challenging approach that has not been explored until now. The objectives in BiSSL are connected by substituting the lower-level backbone solution for the upper-level backbone parameters, while simultaneously enforcing the lower-level backbone solution to resemble the upper-level backbone parameters. This approach more explicitly captures the interdependence between pretext pre-training and downstream fine-tuning, potentially leading to a lower-level backbone better aligned with the downstream task. Figure 1 compares the conventional SSL pipeline with our suggested pipeline involving BiSSL. Additionally, we propose a training algorithm for BiSSL and demonstrate that it consistently improves or maintains comparable downstream performance across a range of image classification datasets. For our experiments, we use SimCLR (Chen et al., 2020b) to pre-train a ResNet-18 backbone (He et al., 2016) on the unlabeled partition of the STL10 dataset (Coates et al., 2011), a setup offering suitable model capacity and dataset complexity while being less resource-intensive than larger-scale alternatives. The code implementation and pre-trained model weights are publicly available. [^1] https://github.com/ICLR25-10484/ICLR25_10484_BiSSL
@@ 3.3 Upper-level Derivative
Due to the dependence of the lower-level solution on the upper-level parameters, the first term of equation 4 includes the implicit gradient (IG) of the implicit function ${\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D})$. To simplify notation, we let $\nabla_{{\bf\it\xi}}h({\bf\it\xi})|_{{\bf\it\xi}={\bf\it\psi}}:=\nabla_{{\bf\it\xi}}h({\bf\it\psi})$ when it is clear from context which variables are differentiated with respect to. Following an approach similar to Rajeswaran et al. (2019), with details on the derivations and underlying assumptions outlined in Section A.1 of Appendix A, the IG in equation 4 can be explicitly expressed as
@@ 3.4 Training Algorithm and Pipeline
From Section A.1 in Appendix A, we get that ${\bf\it\theta}^{\ast}_{P}({\bf\it\theta}_{D})$ must fulfill the stationary condition $\nabla_{{\bf\it\theta}}\quantity(\mathcal{L}^{P}({\bf\it\theta},{\bf\it\phi}_{P})+\lambda r({\bf\it\theta}_{D},{\bf\it\theta}))|_{{\bf\it\theta}={\bf\it\theta}^{\ast}_{P}({\bf\it\theta}_{D})}=\mathbf{0}$ to justify the explicit expression of the implicit gradient in equation 6. This means that executing Algorithm 1 using random initializations of ${\bf\it\theta}$ and ${\bf\it\phi}_{P}$ will likely not suffice. The same applies to ${\bf\it\phi}_{D}$, as a random initialization of ${\bf\it\phi}_{D}$ typically leads to rapid initial changes of the backbone parameters ${\bf\it\theta}_{D}$ during fine-tuning. This would then likely violate the assumed stationary condition due to the dependence between ${\bf\it\theta}_{D}$ and ${\bf\it\theta}_{P}$ through the regularization objective $r$. Figure 1 illustrates the suggested pipeline alongside the conventional SSL pipeline. First conventional pretext pre-training is performed on the unlabeled dataset $\mathcal{D}^{P}$ to obtain initializations of ${\bf\it\theta}$ and ${\bf\it\phi}_{P}$. Next, the downstream head is fitted on top of the frozen backbone ${\bf\it\theta}$ using the downstream dataset $\mathcal{D}^{D}$, which provides an initialization of the downstream head parameters ${\bf\it\phi}_{D}$. Then, BiSSL training is conducted as outlined in Algorithm 1, yielding an updated configuration of backbone parameters ${\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D})$. These updated backbone parameters are subsequently used as an initialization for the final supervised fine-tuning on the downstream task.
@@ 4.2.2 BiSSL Setup
The backbone ${\bf\it\theta}$ and projection head ${\bf\it\phi}_{P}$ are initialized by self-supervised pre-training using a setup almost identical to the baseline pretext task training setup in Section 4.2.1. The only difference is that this training stage is conducted for $500$ epochs instead of $600$ epochs, and that the peak base learning rate is set to $1.0$ instead of $4.8$. This adjustment is made because the BiSSL training stage will conduct what is roughly equivalent to $100$ pretext epochs, as detailed more specifically in the composite configuration paragraph below. This ensures that the total number of pretext pre-training steps is comparable to those conducted in the baseline setup.
@@
STL10L $90.2\pm 0.1$ $90.3\pm 0.1$ $-0.1$ $\mathbf{99.7\pm 0.0}$ $99.6\pm 0.0$ $\mathbf{+0.1}$
Flowers $\mathbf{74.8\pm 0.2}$ $73.4\pm 0.4$ $\mathbf{+1.4}$ $89.8\pm 0.3$ $90.0\pm 0.4$ $-0.2$
Cars $73.0\pm 0.4$ $72.7\pm 0.5$ $+0.3$ $91.5\pm 0.3$ $91.4\pm 0.4$ $+0.1$
Aircrafts $46.9\pm 0.5$ $46.1\pm 0.9$ $+0.8$ $78.9\pm 0.4$ $79.3\pm 0.6$ $-0.4$
DTD $\mathbf{51.8\pm 0.5}$ $49.3\pm 0.5$ $\mathbf{+2.5}$ $\mathbf{79.9\pm 0.3}$ $79.1\pm 0.4$ $\mathbf{+0.8}$
Pets $\mathbf{67.8\pm 0.2}$ $65.0\pm 0.5$ $\mathbf{+2.8}$ $\mathbf{92.3\pm 0.3}$ $90.7\pm 0.3$ $\mathbf{+1.6}$
FMNIST $94.3\pm 0.2$ $94.1\pm 0.1$ $+0.2$ $100.0\pm 0.0$ $100.0\pm 0.0$ $0.0$
CIFAR10 $93.9\pm 0.1$ $93.8\pm 0.1$ $+0.1$ $\mathbf{99.9\pm 0.0}$ $99.8\pm 0.0$ $\mathbf{+0.1}$
CIFAR100 $73.0\pm 0.1$ $73.2\pm 0.2$ $-0.2$ $\mathbf{93.7\pm 0.1}$ $92.8\pm 0.1$ $\mathbf{+0.9}$
Caltech-101 $\mathbf{80.6\pm 0.7}$ $78.1\pm 0.5$ $\mathbf{+2.5}$ $\mathbf{95.5\pm 0.2}$ $94.7\pm 0.2$ $\mathbf{+0.8}$
Food $72.0\pm 0.2$ $71.7\pm 0.2$ $+0.3$ $90.4\pm 0.1$ $90.4\pm 0.1$ $0.0$
SUN397 $\mathbf{41.1\pm 0.2}$ $40.0\pm 0.3$ $\mathbf{+1.1}$ $\mathbf{71.0\pm 0.2}$ $69.9\pm 0.4$ $\mathbf{+1.1}$
CUB200 $\mathbf{47.1\pm 0.4}$ $45.7\pm 0.4$ $\mathbf{+1.4}$ $\mathbf{72.1\pm 0.3}$ $70.7\pm 0.6$ $\mathbf{+1.4}$
VOC07 $*\mathbf{60.4\pm 0.1}$ $*58.6\pm 0.3$ $\mathbf{+1.8}$- -- --
@@ 4.3.1 Performance over varying pre-training epochs
Table 1 demonstrates that BiSSL can significantly enhance classification accuracies across various datasets. To determine whether this improvement persists across different levels of pretext pre-training, additional experiments are conducted using varying numbers of total pre-training epochs. The accuracies are compared with the total number of pretext pre-training epochs kept roughly the same, so the pretext warm-up in the BiSSL pipeline is conducted for $100$ epochs less than the baseline pretext pre-training, as justified in the “Composite Configuration Details of BiSSL” paragraph of Section 4.2.2. The flowers dataset, which showed substantial benefits from BiSSL on the top-1 classification accuracies, is selected for these experiments. The results presented in Figure 2 reveal that BiSSL generally sustains the relative performance gap over the baseline, regardless of the pre-training duration. This suggests that the advantages conferred by BiSSL are not contingent on the amount of pre-training. Rather, this indicates that BiSSL may provide a more efficient learning trajectory, stemming from the enhanced information sharing it facilitates between the pretext and downstream tasks.
@@ 4.4 Visual Inspection of Latent Features
Figure 3: Visualization of features from backbones trained using pretext pre-training exclusively and backbones derived from lower-level backbones obtained after applying BiSSL, respectively. Features are extracted from the test partition of the flowers dataset. Each color represents a different class. Details are outlined in Section C.1 of Appendix C
@@ 5.1 Future Work
Formulating the self-supervised pipeline as a bilevel optimization problem offers various strategies with trade-offs in computational complexity and theoretical justification. While this study presents a promising approach for improving downstream performance, further investigation of alternative formulations is needed to identify setups are are potentially more optimal. Although BiSSL is theoretically applicable to any downstream task and model size, our experiments focused on small-scale image classification due to resource constraints. Therefore, it remains uncertain whether BiSSL can scale to larger setups and tasks. Additionally, a potential future advancement would integrating more novel methods for solving BLO problems, which promise benefits in terms of reduced computational costs and improved solution convergence (Zhang et al., 2023a; Yang et al., 2021; Choe et al., 2023; Huang, 2024). Lastly, the current BiSSL framework relies on full access to pre-training data and pretext tasks. Future research could investigate the use of only a subset of pre-training data and alternative pretext tasks to maintain BiSSL’s benefits under these conditions.
@@ Acknowledgments
The authors would like to thank Sijia Liu and Yihua Zhang (Michigan State University) providing valuable feedback in a discussion, which helped to refine and solidify our perspective on the topic of integrating bilevel optimization in deep learning.
Appendix A Theoretical Insights and Framework Comparisons in BiSSL
A.1 Derivation of the Implicit Gradient
@@
A.2 Distinction from Bilevel Optimization in Meta-Learning
While bilevel optimization (BLO) has been applied in meta-learning frameworks such as MAML (Finn et al., 2017), Sign-MAML (Fan et al., 2021) and iMAML (Rajeswaran et al., 2019), BiSSL represents a distinct application and implementation of BLO, tailored for the challenges of self-supervised learning (SSL). In the aforementioned works, BLO is primarily utilized to address few-shot learning scenarios, focusing on efficiently adapting models to new tasks with minimal data. Conversely, BiSSL applies BLO to concurrently manage the more complex task of self-supervised pretext pre-training with downstream fine-tuning. Another key distinction is that in meta-learning, the upper- and lower-level objectives are closely related, with the upper-level objective formulated as a summation of the lower-level tasks. In contrast, BiSSL involves fundamentally distinct objectives at each level, utilizing separate datasets and tasks for pre-training and fine-tuning. This design allows BiSSL to better align the pre-trained model with the requirements of a specific downstream task. Conversely, the BLO in meta-learning aims to broadly generalize across a wide range of tasks, prioritizing adaptability rather than task-specific optimization. Additionally, unlike BiSSL, the meta-learning frameworks discussed reinitialize the lower-level backbone parameters with a copy of the upper-level parameters at every iteration. In BiSSL, the closest comparable mechanism is the occasional update of the upper-level backbone using an EMA update with the lower-level parameters (see Algorithm 1), though this occurs far less frequently. Lastly, while the meta-learning with BLO approaches often target smaller models and datasets, BiSSL is aimed for larger-scale setups characteristic of modern SSL workflows.
@@ B.5 Downstream Fine-Tuning after BiSSL
STL10L $0.0005$ $0.00043$
Flowers $0.0009$ $0.00005$
Cars $0.0293$ $0.00851$
Aircrafts $0.011$ $0.00612$
DTD $0.0008$ $0.00764$
Pets $0.0002$ $0.00543$
FashionMNIST $0.0022$ $0.00876$
CIFAR10 $0.0003$ $0.00991$
CIFAR100 $0.0012$ $0.00422$
Caltech-101 $0.0008$ $0.00011$
Food $0.006$ $0.0095$
SUN397 $0.0011$ $0.00028$
CUB200 $0.035$ $0.00868$
VOC2007 $0.0002$ $0.00015$